OpenStudy (anonymous):

Write an expression to describe a rule for the sequence. Then find the 100th term. -4, 1, 6, 11, 16, . . .

OpenStudy (anonymous):

it always adds 5

OpenStudy (anonymous):

oh yeah.. then the 100th term is 16 times 95 which is 1520

OpenStudy (anonymous):

i think...

OpenStudy (anonymous):

I am not familiar with sequences. how did you get 16, and 95?

OpenStudy (anonymous):

i have no idea... now that i think of it i really dont know -_-

OpenStudy (anonymous):

because there are 5 so far, and we want #100, so it is 95 left right? and the last one is 16 :)

OpenStudy (anonymous):

is -4 for n=1 ?

OpenStudy (anonymous):

do we start the sequence with "1"?

OpenStudy (anonymous):

first term - term n=1 - is -4, right ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok, then this formula will work: multiply n by 5 and subtract 9 if n=1, we have n*5=5-9=-4 if n=2, we have 2*5=10-9=1 if n=3, we have 3*5=15-9=6

OpenStudy (anonymous):

\[n_{th} term=5 \times n -9\]

OpenStudy (anonymous):

you lost me there dude.. i gonna be honest with u, i have no idea what your saying :)

OpenStudy (anonymous):

I tried some formulas and came up with the one above :) in the post, I just shared with you how I checked it look at the equation above, try it for some term of the sequence

OpenStudy (john_es):

\[a_n=a_0+d(n-1)\\ d=5\\ a_0=-4\\ a_n=-4+5(n-1)=5n-9\]

OpenStudy (john_es):

\[a_{100}=5\cdot100-9=491\]